Find the eigenvalues and eigenvectors of the matrices

1) [ 9 3 ]

2 9


2) 2 0 1

[ 0 2 0 ]

1 0 2


1
Expert's answer
2021-05-26T17:54:22-0400

(1) "\begin{bmatrix}
 9 & 3 \\
 2 & 9
\end{bmatrix}"

Start from forming a new matrix by subtracting λ

 from the diagonal entries of the given matrix: 

"\left[\begin{array}{cc}9 - \lambda & 3\\2 & 9 - \lambda\end{array}\right]"

The determinant of the obtained matrix is "\lambda^{2} - 18 \lambda + 75"

Solve the equation

The roots are comes to be "\lambda_{1} = 9 - \sqrt{6}" , "\lambda_{2} = \sqrt{6} + 9"

These are the eigenvalues.


Next, find the eigenvectors.

a. "\lambda = 9 - \sqrt{6}"

"\left[\begin{array}{cc}9 - \lambda & 3\\2 & 9 - \lambda\end{array}\right] = \left[\begin{array}{cc}\sqrt{6} & 3\\2 & \sqrt{6}\end{array}\right]"

The null space of this matrix is"\left\{\left[\begin{array}{c}- \frac{\sqrt{6}}{2}\\1\end{array}\right]\right\}"

This is the eigenvector.



b. "\lambda = \sqrt{6} + 9"

"\left[\begin{array}{cc}9 - \lambda & 3\\2 & 9 - \lambda\end{array}\right] = \left[\begin{array}{cc}- \sqrt{6} & 3\\2 & - \sqrt{6}\end{array}\right]"

The null space of this matrix is "\left\{\left[\begin{array}{c}\frac{\sqrt{6}}{2}\\1\end{array}\right]\right\}"

This is the eigenvector.


(2)

"\begin{bmatrix}
 2 & 0 & 1 \\
 0& 2 & 0 \\
1 & 0 & 2
\end{bmatrix}"

as above done we well proceed the same

Start from forming a new matrix by subtracting λ

λ from the diagonal entries of the given matrix: 

"\left[\begin{array}{ccc}2 - \lambda & 0 & 1\\0 & 2 - \lambda & 0\\1 & 0 & 2 - \lambda\end{array}\right]."


The determinant of the obtained matrix is 

"- \lambda^{3} + 6 \lambda^{2} - 11 \lambda + 6"

Solve the equation "\lambda_{1} = 3" , "\lambda_{2} = 2" , "\lambda_{3} = 1"

These are the eigenvalues.


Next, find the eigenvectors.

a. "\lambda = 3"

"\left[\begin{array}{ccc}2 - \lambda & 0 & 1\\0 & 2 - \lambda & 0\\1 & 0 & 2 - \lambda\end{array}\right] = \left[\begin{array}{ccc}-1 & 0 & 1\\0 & -1 & 0\\1 & 0 & -1\end{array}\right]"


The null space of this matrix is "\left\{\left[\begin{array}{c}1\\0\\1\end{array}\right]\right\}"


This is the eigenvector.


b. "\lambda = 2"

"\left[\begin{array}{ccc}2 - \lambda & 0 & 1\\0 & 2 - \lambda & 0\\1 & 0 & 2 - \lambda\end{array}\right] = \left[\begin{array}{ccc}0 & 0 & 1\\0 & 0 & 0\\1 & 0 & 0\end{array}\right]"

The null space of this matrix is"\left\{\left[\begin{array}{c}0\\1\\0\end{array}\right]\right\}"

This is the eigenvector.


c. "\lambda = 1"

"\left[\begin{array}{ccc}2 - \lambda & 0 & 1\\0 & 2 - \lambda & 0\\1 & 0 & 2 - \lambda\end{array}\right] = \left[\begin{array}{ccc}1 & 0 & 1\\0 & 1 & 0\\1 & 0 & 1\end{array}\right]"

The null space of this matrix is"\left\{\left[\begin{array}{c}-1\\0\\1\end{array}\right]\right\}"

This is the eigenvector.